What substance can gain 40.0 KJ and only gain 25 C if 1.616 kg of the substance is present?
1
Expert's answer
2020-06-01T14:24:46-0400

Q = mct

C = Q/mt = 40×103J/1.616kg×25°C = 990.1 J/kg×°C

It is the specific heat capacity of air.